LIF SPECTROSCOPY OF A CARBON CHAIN RADICAL PRODUCED BY A HC_{3}N DISCHARGE

HIDEYUKI TAKADA, \underline{YOSHIHIRO SUMIYOSHI}, and YASUKI ENDO, Department of Pure and Applied Sciences, College of Arts and Sciences, The University of Tokyo, 153-8902 Tokyo, Japan; KENNOSUKE HOSHINA, Department of Chemistry, School of Science, The University of Tokyo, 113-0033 Tokyo, Japan.

New electronic transitions of a carbon chain radical containing a nitrogen atom have been observed by LIF spectroscopy. The radical was produced in a supersonic jet by a pulsed discharge of HC_{3}N diluted to 0.3% with 3atm Ar. We confirmed that the radical contains a nitrogen atom from the fact that the spectrum was not observed when C_{2}H_{2} was used as a precursor. About ten vibronic bands with a 50nsec radiative lifetime have been observed in a region from 27000 to 27400 cm^{−1}. The observed bands could be classified as ^{2}Π_{3/2}-^{2}Π_{3/2} and ^{2}Σ −^{2} Σ vibronic bands. The lower state rotational constants of the vibronic bands have been determined to be 0.07537(1) and 0.07932(5) cm^{−1} for the Π and Σ states, respectively. The upper states of ^{2}Σ −^{2} Σ bands have very large spin-rotation constants, which should be denoted as Σ^{(±)} subjected to a small Renner-Teller interaction. The new molecule has been tentatively assigned to be the C_{4}N radical. We performed an UMP2 level ab initio calculation with 6-311G** basis sets for the C_{4}N radical. The estimated rotational constant in the ground state was 0.077 cm^{−1}. This value agrees well with the observed rotational constants.